Hello please make sure for ever step possible ok?

(f) (f) (a) Ignore friction and determine an expression for the distance d the boxes travel before coming to rest. Choose the floor as the position for zero gravitational potential energy. (Use the following as necessary: m m2, v, and g for the acceleration due to gravity.) m1 +m2 V 2m (b) Is the work done on box 2 by the rope positive, negative, or zero? positive negative zero (c) Determine an expression for the work done on box 1 by the rope. (Use the following as necessary: m m2, v, and g for the acceleration due to gravity.)
Answer #1

(a)

From law of energy conservation,

(1/2)*m1*v^2 + (1/2)*m2*v^2 = m2*g*d

(m1 + m2)*v^2 = 2*m2*g*d

d = (m1 + m2)*v^2 / 2*m2*g

(b)

Work done on box 2 by the rope is positive.

(c)

Work done on box 1 by the rope = KE of box 1

W = m1*v^2 / 2